# The LHAASO Collaboration

### v2

@@ Appendix A LHAASO detectors
The data with energy above 25 TeV collected by KM2A are used in this analysis. KM2A comprises 5,216 electromagnetic particle detectors (EDs) and 1,188 underground muon detectors (MDs) spread over an area of approximately 1.3 km2 [35]. The KM2A was incrementally developed and began operations, with half of the array starting scientific observations at the end of 2019, three-quarters operational by December 2020, and full operations officially launched on July 31, 2024, with a total exposure time of 1570.42 days. The ratio of measured muons to electrons is utilized to discriminate primary $\gamma$-rays from cosmic nuclei. Energies are estimated based on particle density at a radius of 50 m (denoted as $\rho_{50}$) across various zenith angles [19]. The energy resolution is better than 20% for energies above 100 TeV within the zenith angle range of $0^{\circ}-35^{\circ}$. The data sets are categorized into five groups per decade based on reconstructed energy. The sky map is represented in celestial coordinates, divided into a grid of $0.1^{\circ}\times 0.1^{\circ}$ pixels, with each pixel filled with the number of detected events corresponding to their reconstructed arrival direction. The “direct integration method” is employed to estimate the cosmic ray background. To optimize array performance, several event selection criteria are implemented. The data selection criteria used in this study align with the guidelines established in the publication by the collaboration in the Chinese Physics C (CPC) journal in 2021 [19].
The VHE data ($0.8\ {\rm TeV}<E<31\ {\rm TeV}$ ) were collected by the whole WCDA. The data used in this study were obtained from the full-array operation of the WCDA from March 5, 2021, to July 31, 2024, with a total exposure time of 1124 days. The zenith angle of the shower was required to be less than 50 degrees. The Gamma/Proton separation parameter $\mathcal{P}_{\rm{incness}}$ [36], is required to be less than 1.12, 1.02, 0.90, 0.88, 0.88, 0.84, and 0.84 for segments with $N_{\rm{hit}}$ value of [3060),[60100), [100200), [200300), [300500), [500800) and [8002000]. For each $N_{\rm{hit}}$ segment, the sky map in celestial coordinates (right ascension and declination) was divided into a grid of $0.1^{\circ}\times 0.1^{\circ}$, which was filled with the number of the detected events according to their reconstructed arrival direction. The “direct integration method” [37] is adopted to estimate the number of CR background events in each pixel.
Table 1: Energy bin definitions for WCDA (upper) and KM2A (lower) and the characteristics of LS I +61 303 at each energy bin. Note that the definitions for WCDA and KM2A are different: WCDA data is binned depending on the number of triggered detector units($N_{\rm{hit}}$), and KM2A data is divided by the reconstructed energy.
@@ Appendix B Data analysis
Considering the influence from Galactic diffuse $\gamma$-ray emission (GDE) and nearby bright sources on LS I +61 303, we define a 4-radius circular ROI around LS I +61 303 , centered at R.A. = 40.1, Decl. = 61.2 (see Fig. 4). The only significant source inside this region is J0248+6021, which lies close to LS I +61 303.
Figure 4: Observed maps within ROI from WCDA (A) and KM2A (B).
@@
Figure 5 shows the residual map after subtracting all sources within the ROI. No significant excess is visible in this map. The distribution of TS values is converted into statistical significance using Wilks’s theorem and is well fit by a Gaussian. For WCDA the fit yields a mean of $0.055\pm 0.010$ and a standard deviation of $0.978\pm 0.006$. For KM2A we obtain a mean of $0.033\pm 0.007$ and a standard deviation of $0.996\pm 0.005$.
Figure 5: 2D residual map from WCDA (A) and KM2A (B).
To evaluate the pointing error at the declination of LS I +61 303, we’ve analysed Crab’s position using events with zenith angle >>30 and adjusted the distribution of zenith angle to the path of LS I +61 303. For KM2A, the largest deviation is $0.036^{\circ}\pm 0.015^{\circ}$ in RA and $0.028^{\circ}\pm 0.016^{\circ}$ in DEC. The pointing error of WCDA is calibrated using extra-galactic source 1ES 1959+650, which is at similar declination of LS I +61 303. The deviation in RA and DEC are $0.029\pm+0.062$ and $0.009\pm 0.024$. It should be noted that the deviation is dominated by statistics, and the pointing accuracy is calibrated to be better than $0.03^{\circ}$ using all the events from Crab.

### v3

@@ Appendix A LHAASO detectors
The VHE data ($0.8\ {\rm TeV}<E<31\ {\rm TeV}$ ) were collected by the whole WCDA. The data used in this study were obtained from the full-array operation of the WCDA from March 5, 2021, to July 31, 2024, with a total exposure time of 1124 days. The zenith angle of the shower was required to be less than 50 degrees. The $\gamma$/Proton separation parameter $\mathcal{P}$ (a parameter that characterizes the ‘clumpiness’ of the air shower, termed PINCness, is detailed in [3]) is required to be less than 1.12, 1.02, 0.90, 0.88, 0.88, 0.84, and 0.84 for segments with $N_{\rm{hit}}$ value of [3060), [60100), [100200), [200300), [300500), [500800) and [8002000]. For each $N_{\rm{hit}}$ segment, the sky map in celestial coordinates (right ascension and declination) was divided into a grid of $0.1^{\circ}\times 0.1^{\circ}$, which was filled with the number of the detected events according to their reconstructed arrival direction. The ‘direct integration method’ is adopted to estimate the number of CR background events in each pixel.
The data with energy above 25 TeV collected by KM2A are used in this analysis. KM2A comprises 5,216 electromagnetic particle detectors (EDs) and 1,188 underground muon detectors (MDs) spread over an area of approximately 1.3 km2 [28]. The KM2A was incrementally developed and began operations, with half of the array starting scientific observations at the end of 2019, three-quarters operational by December 2020, and full operations officially launched on July 31, 2024, with a total exposure time of 1570.42 days. Only showers with reconstructed zenith angle less than 50 are used. The ratio of measured muons to electrons is utilized to discriminate primary $\gamma$-rays from cosmic nuclei. Energies are estimated based on particle density at a radius of 50 m (denoted as $\rho_{50}$) across various zenith angles. The energy resolution is better than 20% for energies above 100 TeV within the zenith angle range of $0^{\circ}$$35^{\circ}$. The data sets are categorized into five groups per decade based on reconstructed energy. The sky map is represented in celestial coordinates, binned into a grid of $0.1^{\circ}\times 0.1^{\circ}$ pixels, with each pixel filled with the number of detected events corresponding to their reconstructed arrival direction. The ‘direct integration method’ is employed to estimate the cosmic ray background. To optimize array performance, several event selection criteria are implemented. The data selection criteria published in the KM2A performance paper[6] are used in this study.
Table S1: Energy bin definitions for WCDA (upper) and KM2A (lower) and the characteristics of LS I +61 303 at each energy bin. Note that the definitions for WCDA and KM2A are different: WCDA data is binned depending on the number of triggered detector units($N_{\rm{hit}}$), and KM2A data is divided by the reconstructed energy.
@@ Appendix B Data analysis
B.1 Likelihood fitting
Considering the influence from Galactic diffuse gamma-ray emission (GDE) and 1LHAASO J0249+6022 close to LS I +61 303, we define a 4-radius circular ROI around LS I +61 303, centered at R.A. = 40.1, Decl. = 61.2 (see Fig. S1). The data analysis is based on a 3-dimensional likelihood fitting method, where the position and spectrum for each source in the Region of Interest (ROI) are fitted simultaneously. The morphology of the source is described by a 2D Gaussian function. Two kinds of spectrum forms are tested in this analysis, which are the simple power-law function and the exponential cut-off function:
$dN/dE=F\times(E/E_{0})^{-\alpha}$ (1)
$dN/dE=F\times(E/E_{0})^{-\alpha}\times e^{-E/E_{c}}$ (2)
Where $F$ is the differential flux normalization at $E_{0}$. $E_{0}$ is the reference energy, which is chosen to be 50 TeV for KM2A and 3 TeV for WCDA separately. $\alpha$ is the spectral index, and Ec is the cut-off energy.
The observed number of events in each pixel is assumed to follow a Poisson distribution, with the mean value being the sum of cosmic-ray background counts and expected gamma-ray counts from sources and the GDE. The expected gamma-ray counts are calculated by folding the source model through the instrument response (obtained from detector simulations). The likelihood value $\mathcal{L}$ is defined as the product of the probabilities of each pixel. It should be explicitly noted that in this 3D likelihood maximization, the cosmic-ray background is fixed to the value derived from the direct integration method. This approach corresponds to the Cash statistic [18]. Ignoring the statistical uncertainty of the background in the fit is practically justified because the direct integration method leverages a massive off-source data pool (resulting in an effective ON/OFF exposure ratio $\alpha\ll 1$), rendering the statistical error on the background estimation negligible compared to the fluctuations of source counts. However, in scenarios where the cosmic-ray background heavily dominates, its associated uncertainty becomes critical and should be explicitly evaluated (e.g., [13]). Conversely, in the >100>100 TeV regime, contributions from the GDE and 1LHAASO J0249+6022 become negligible. This spatial simplicity naturally allows us to employ the standard Li & Ma formula [35], which rigorously incorporates the background statistical uncertainty but is otherwise difficult to apply in complex, multi-component regions.
The likelihood ratio test is used to perform the hypothesis testing. The Test Statistic (TS) is defined as:
$\rm TS=-2(ln\mathcal{L}_{0}-ln\mathcal{L}_{1})$ (3)
@@ B.1 Likelihood fitting
Figure S1: Test statistic (TS) maps of the region around LS I +61 303 observed by WCDA (A) and KM2A (B). The color scale represents the TS value of the excess. The target source LS I +61 303 is clearly visible in the center. The neighboring source 1LHAASO J0249+6022, located $\sim 1.9^{\circ}$ away, is likely a PWN or pulsar halo associated with PSR J0248+6021. The large-scale extended features are attributed to the Galactic diffuse emission (GDE).
B.2 Results
[…]
@@ B.2 Results
Fig. S2 shows the residual map after subtracting all modeled sources within the ROI, in which no significant excess is visible. The distribution of TS values from the map is converted into statistical significance according to Wilks’ theorem and is well-fitted by a Gaussian distribution, as shown in Fig. S3. For WCDA, the fit yields a mean of $0.05\pm 0.02$ and a standard deviation of $1.06\pm 0.02$. For KM2A, we obtain a mean of $0.03\pm 0.02$ and a standard deviation of $0.97\pm 0.02$.
Figure S2: The 2D residual test statistic (TS) maps from WCDA (A) and KM2A (B) are shown after subtracting the best-fit models for LS I +61 303, 1LHAASO J0249+6022, and the Galactic diffuse emission. Each pixel represents the TS value derived from a likelihood ratio test. No residual excess with $\sqrt{\text{TS}}>3$ is observed in either map.
Figure S3: The 1D histograms show the distribution of $\sqrt{\text{TS}}$ values extracted from the residual maps (Fig. S2) for WCDA (A) and KM2A (B). The red solid lines represent the best-fit results using a Gaussian function. Both distributions are consistent with a standard Gaussian (mean $\approx 0$, $\sigma\approx 1$), indicating that the residuals are dominated by statistical fluctuations and the background is properly modeled.
To evaluate the systematic pointing uncertainty at the declination of LS I +61 303, we analyzed the position of the Crab Nebula using events with zenith angles $\theta>30^{\circ}$, weighting the zenith distribution to match the transit path of LS I +61 303. For KM2A, the maximum observed deviations were $0.036^{\circ}\pm 0.015^{\circ}$ in R.A. and $0.028^{\circ}\pm 0.016^{\circ}$ in Decl. Although these values are likely dominated by statistical limitations, we conservatively adopt a systematic pointing error of $0.05^{\circ}$ for KM2A. For WCDA, the same procedure yielded a positional deviation of less than $0.01^{\circ}$. We further cross-verified this accuracy using multiple extragalactic sources, including Mrk 421, Mrk 501, and 1ES 1959+650 (located at a declination similar to LS I +61 303); these cross-checks confirm that the WCDA pointing error is better than $0.03^{\circ}$.
Figure S4: Orbital modulations at multiple wavelengths. (A,B,C) are the same with Fig. 3. (D) shows the light curve derived from simultaneous Fermi-LAT observations. (E) Although the Swift-XRT [24] and VERITAS [29] data were not taken contemporaneously with this work, their peak fluxes both occur near $\phi\sim 0.6$.
B.3 Fermi-LAT data analysis
To obtain the Fermi orbital light curve contemporaneous with LHAASO observations (since the source exhibits different orbital modulation patterns at different superorbital phases), we analyzed approximately three years of data (MJD 59415–60523). The PASS 8 SOURCE class events were extracted from a $20^{\circ}\times 20^{\circ}$ region of interest (ROI) centered on LS I +61 303, with an energy range of 0.1–300 GeV. The analysis was performed using Fermitools v2.2.0, adopting the instrument response function P8R3_SOURCE_V3, and the background model was constructed based on the 4FGL-DR4 catalog [2], along with the Galactic diffuse component (gll_iem_v07.fits) and the isotropic extragalactic component (iso_P8R3_SOURCE_V3_v1.txt). The globally fitted model was subsequently used as the baseline for the orbital flux reconstruction. In practice, the orbital period was evenly divided into 16 phase bins, after which we performed phase-resolved likelihood analyses. Specifically, events were re-filtered within each phase interval following period folding, and the likelihood fit was independently repeated to derive the flux in each phase bin. Given the smoothing effect introduced by folding over the orbital period of LS I +61 303, we assumed that background sources do not vary with this orbital phase. Therefore, during the phase-resolved fitting, only the spectral parameters of LS I +61 303 and normalizations of the diffuse components were left free, while all other background sources were fixed to their global best-fit values obtained above.
Appendix C Binary parameter setting
In the discussion, we adopt parameter values established in previous studies as the baseline. The temperature of the Be star is taken to be 22.5 kK [38], corresponding to a characteristic blackbody photon energy of $E\sim k_{\rm B}T=1.94$ eV. The peak of the SED occurs at $E_{\rm peak}\approx 5.5$ eV, which lies in the ultraviolet band. According to the model proposed by Millar & Marlborough (1998) [40], the temperature in the inner region of the circumstellar disk is approximately 60% of the stellar temperature, yielding $T_{\rm disk,0}\approx 0.6T_{*}\approx 13500$ K. For viscous-disk, the radial temperature profile can be estimated using $T(r)=T_{\rm disk,0}(r/R_{*})^{-3/4}$ [36]. To determine the specific environment encountered by the compact object, we must first constrain the orbital distance. Since the orbital eccentricity of the binary is known ($e\sim 0.54$ [11]), the binary separation at periastron can be calculated as $r_{\rm peri}=a(1-e)$, where the semi-major axis $a=0.42$ AU is obtained from Kepler’s law. This gives $r_{\rm peri}\approx 0.19$ AU. Under additional parameter assumptions ($R_{*}=10R_{\odot}$ and a disk radius $R_{d}=7R_{*}\approx 0.33$ AU [49]), it becomes evident that the compact object may undergo a disk-crossing event when passing through periastron.
To further estimate the matter density variations encountered by the compact object during the potential periastron disk-crossing event, here we adopt canonical values for the mass-loss rate $\dot{M}\sim 5\times 10^{-8}M_{\odot}~{\rm yr}^{-1}$ and velocity $v_{w}=10^{8}~{\rm cm}~{\rm s}^{-1}$. At periastron, this yields $n(r_{\rm p})\approx 1.9\times 10^{8}~{\rm cm}^{-3}$. Within the Be-star circumstellar disk, the particle density is expected to exceed that of the stellar wind by approximately 2–4 orders of magnitude, depending on the disk structure and radial distance (e.g., [42]). For the purposes of our estimates, we adopt a fiducial disk density of $n_{\rm disk,peri}\sim 10^{11}~{\rm cm}^{-3}$ at periastron. In the $pp$ scenario, the interaction efficiency can be estimated as $f_{pp}=t_{\rm esc}/(t_{pp}+t_{\rm esc})$, where the escape/advection timescale is aproximated by $t_{\rm esc}\sim l/v$. For a compact interaction region with a characteristic size $l\sim 10^{12}$ cm and a flow velocity $v\sim 10^{9}~{\rm cm}~{\rm s}^{-1}$, this yields $t_{\rm esc}\sim 10^{3}$ s. As a result, we obtain $f_{pp}\approx 0.06$ during disk-crossing episodes, while it decreases to $f_{pp}\approx 10^{-4}$ outside these phases.
The KM2A energy flux in the 25–100 TeV band is measured to be $F_{\gamma,{\rm 25-100TeV}}=1.05\times 10^{-13}~{\rm erg}~{\rm cm}^{-2}~{\rm s}^{-1}$. Assuming a source distance of $d\approx 2.0$ kpc, the corresponding gamma-ray luminosity is $L_{\gamma}=4\pi d^{2}F_{\gamma}\approx 5\times 10^{31}~{\rm erg}~{\rm s}^{-1}$. The proton power required to sustain this gamma-ray luminosity is given by $L_{p}\approx 3L_{\gamma}/f_{pp}$, where $f_{pp}$ denotes the effective $pp$ interaction efficiency. To assess whether such a proton power can be supplied by a pulsar, we estimate the spin-down luminosity assuming a neutron star with period $P\simeq 270$ ms [47], moment of inertia $I\approx 10^{45}~{\rm g}~{\rm cm}^{2}$, and period derivative $\dot{P}=3\times 10^{-15}~{\rm s}~{\rm s}^{-1}$ [37]. The spin-down power is then $\dot{E}=4\pi^{2}I\dot{P}/P^{3}\approx 6\times 10^{33}~{\rm erg}~{\rm s}^{-1}$. However, integrating the measured soft power-law spectrum ($\Gamma\approx 3.0$) above 1 TeV yields a bolometric gamma-ray luminosity of $L_{\gamma,>1{\rm TeV}}\approx 1.7\times 10^{33}~{\rm erg}~{\rm s}^{-1}$. In a purely hadronic scenario with an interaction efficiency of $f_{pp}\approx 0.06$, the required proton power would be $L_{p}\approx 3L_{\gamma}/f_{pp}\approx 8.5\times 10^{34}~{\rm erg}~{\rm s}^{-1}$, which exceeds the available pulsar spin-down power ($\dot{E}\approx 6\times 10^{33}~{\rm erg}~{\rm s}^{-1}$) by over an order of magnitude. This energy budget crisis is naturally resolved in a mixed lepto-hadronic scenario. Since electrons are in the fast-cooling regime ($f_{\rm IC}\approx 1$; $t_{IC}\ll t_{\rm esc}$) due to the intense stellar radiation field, they can efficiently power the bulk emission below tens of TeV (we take 25 TeV here, which requires $L_{e}\approx 1.7\times 10^{33}~{\rm erg}~{\rm s}^{-1}$). Protons are thus only required to sustain the high-energy tail (>25>25 TeV, $L_{\gamma}\approx 5\times 10^{31}~{\rm erg}~{\rm s}^{-1}$); with $f_{pp}\approx 0.06$, this demands $L_{p}\approx 2.5\times 10^{33}~{\rm erg}~{\rm s}^{-1}$. The total injection power $L_{\rm inj}\approx 4.2\times 10^{33}~{\rm erg}~{\rm s}^{-1}$ fits comfortably within $\dot{E}$, favoring a mixed origin where protons dominate only the highest energies. The temporal correlation between X-ray and VHE gamma-ray fluxes peaking at $\phi\sim 0.6$ (Fig. S4) implies a common leptonic origin, where the same electron population produces synchrotron X-rays and Inverse Compton gamma-rays. Identifying this distinct leptonic component confirms the validity of the mixed scenario and significantly alleviates the energy budget crisis that arises if the entire broadband emission is attributed solely to hadronic processes.